OpenStudy (daisy.xoxo619):

What interest rate compounded daily(365days/year) is required for a $22000 investment to grow to $38500 in 5 years ? The amount $22,000 will grow to $38500 if the interval rate would be ___% @ganeshie8 @mathmale @mathstudent55

OpenStudy (daisy.xoxo619):

@ganeshie8 @mathmale @mathstudent55

OpenStudy (photon336):

may need to use the formula for compounding but continuously

ganeshie8 (ganeshie8):

For the first part you need to solve r in \[ 38500= 22000*(1+r/(365*100))^{365*5}\]

OpenStudy (wolf1728):

I thought I'd make a graphic to make that equation a little neater. In order to solve the equation for "r", we must take the logarithm of both sides. (See attachment) How do you take the logarithm of (1 +r/36,500) ^ 1,825 ?

OpenStudy (wolf1728):

Okay, I have calculated that formula a little bit more and it is attached. So, I guess the question now becomes what is the logarithm of (1 +r/36,500) ?

OpenStudy (wolf1728):

Okay, let's try this: Attached is a formula from this page: http://www.1728.org/compint3.htm Using the formula exp = log(total / principal) / (compounding periods * years) exp = log(38,500 / 22,000) / (365 * 5) exp = 0.243038048686294 / 1,825 exp = 0.000133171533526737 Now, we need to input that number into this formula: rate = (10^exp -1) * (compounding periods) rate = (10^0.000133171533526737 -1) * (365) rate = (1.00030668580639 -1) * (365) rate = (.00030668580639) * 365 rate = 0.1119403193 rate = 11.194% Which is the daily rate To be extra sure, the calculator here: http://www.1728.org/compint.htm calculates the daily interest rate to be 11.194% Yes, it's just that simple :-)